# LEMUR Neural Network Dataset: Towards Seamless AutoML

### v2

@@ Abstract
Neural networks are fundamental in artificial intelligence, driving progress in computer vision and natural language processing. High-quality datasets are crucial for their development, and there is growing interest in datasets composed of neural networks themselves to support benchmarking, automated machine learning (AutoML), and model analysis. We introduce LEMUR, an open source dataset of neural network models with well-structured code for diverse architectures across tasks such as object detection, image classification, segmentation, and natural language processing. LEMUR is primarily designed to provide a rich source of structured model representations and associated performance data, enabling the fine-tuning of large language models for AutoML applications. Leveraging Python and PyTorch, LEMUR enables seamless extension to new datasets and models while maintaining consistency. It integrates an Optuna-powered framework for evaluation, hyperparameter optimization, statistical analysis, and graphical insights. LEMUR VR extension enables the seamless deployment of models in virtual reality, optimizing their performance on resource-constrained devices. Providing tools for model evaluation, preprocessing, and database management, LEMUR supports researchers and practitioners in developing, testing, and analyzing neural networks. It offers an API that delivers comprehensive information about neural network models and their complete performance statistics with a single request, which can be used in experiments with code-generating large language models. The LEMUR and its plugins are accessible as open source projects under the MIT license at https://github.com/ABrain-One/nn-dataset, https://github.com/ABrain-One/nn-plots and https://github.com/ABrain-One/nn-vr.
@@ 1 Introduction
The rapid advancement of artificial intelligence and machine learning is largely driven by the development of powerful neural network architectures. However, designing, optimizing, and benchmarking these models remains a complex challenge, often requiring extensive manual effort and computational resources. While traditional datasets provide raw data for model training, there is a growing need for structured datasets composed of neural networks themselves, enabling efficient evaluation, benchmarking, and automation in machine learning workflows.
@@
With built-in support for image classification, segmentation, object detection, and natural language processing, LEMUR is a comprehensive resource for research and practical applications. It excels in AutoML by streamlining model selection and hyperparameter tuning while also supporting edge devices for efficient deployment in resource-constrained environments. Large language models, driven by their growing complexity and scale, are increasingly replacing traditional methods across various domains [7, 32], including those focused on the automated exploration of extensive hyperparameter spaces. Our dataset has been utilized by [14, 15], demonstrating its effectiveness in optimizing hyperparameters with large language models.
@@ 1.1 Related Work
In the domain of Neural Networks, datasets are a crucial tool for training, evaluation, and benchmarking processes. Typical datasets in this field include large-scale, labeled datasets such as ImageNet for vision tasks [33] and GLUE for natural language processing [36]. While datasets for images, text, and even biological data have seen extensive development [22, 17, 20, 35, 24], datasets capturing the structures and variations of neural networks themselves have not been curated before.
The motivation behind creating such a dataset came from utilizing PyTorch Vision Library, which provides pre-trained models [29]. Our work expands upon PyTorch’s implementation by normalizing and structuring models into a cohesive dataset that includes architecture details, hyperparameters, and benchmarking metrics. A similar resource is Hugging Face’s robust collection of pre-trained models [38], which, like the PyTorch Vision Library, does not implement normalization that would allow for unified flow and cross-domain evaluations.
Another important consideration is the role of AutoML frameworks, which automate the optimization and evaluation of machine learning models. These frameworks have been shown to support the development of scalable and reproducible workflows by standardizing benchmarking and normalization processes [13, 25]. Furthermore, advances in transfer learning architectures [41] and novel loss functions [21] demonstrate the importance of structured evaluation and maintainability in neural network research. However, these approaches lack integration with cohesive datasets that emphasize the maintainability of neural network structures and workflows. Our work aims to address this gap by incorporating these principles into a unified and extensible dataset designed to support diverse machine learning experiments.
Multiple attempts to create a unified repository of neural networks have been proposed for several different use cases. The CrossedWires dataset [42] examines performance discrepancies between identical models implemented in different frameworks, revealing significant accuracy deviations and therefore underscoring the challenges in achieving cross-framework reproducibility and highlighting the need for standardized evaluation protocols. The Younger dataset [40] pioneers the domain of Artificial Intelligence-Generated Neural Network Architecture (AIGNNA) by compiling 7,629 unique architectures derived from more than 174,000 real-world models in more than 30 tasks, serving as a benchmark for the advancement of graph neural networks and supporting the development of novel neural architectures. The ’deeplearning-models’ repository [26] by Sebastian Raschka is a collection of deep learning architectures and models implemented in TensorFlow and PyTorch, aimed primarily at educational purposes. The Wolfram Neural Net Repository [31] is a public resource hosting an expanding collection of neural network models, suitable for immediate use. The ONNX Model Zoo [27] is a repository of pre-trained, state-of-the-art models in the Open Neural Network Exchange (ONNX) format, that enhances cross-framework compatibility. Collectively, these datasets and repositories enrich the deep learning ecosystem by offering diverse resources tailored to various research and educational needs.
The LEMUR dataset aims to provide an expandable structured repository of standardized neural network implementations accompanied by a unified evaluation and benchmarking framework. The proposed dataset automates performance verification across training hyperparameters and data transformations by generating performance statistics to facilitate the verification of neural network performance. This approach complements existing datasets by offering a foundation for consistent and comparable evaluations across diverse neural network models. By ensuring model consistency and incorporating automated performance tracking, LEMUR addresses reproducibility concerns and facilitates seamless integration with AutoML workflows.
We believe that LEMUR’s focus on standardization and unified evaluation positions it as a valuable asset for advancing reproducibility and efficiency in neural network research and application.
@@ 1.2 Our Contribution
Upon investigating existing datasets designed to collect neural network models and sometimes serve as benchmarks, several critical assumptions were made in the design phase. A reliable neural network dataset should prioritize standardization, extensibility, and versatility. Additionally, an ideal system should facilitate hyperparameter optimization within its mechanism and generate visualizations that summarize training processes.
Standardization ensures that neural network models follow a uniform format, making it easier to compare performance across different architectures. Lack of standardization often leads to discrepancies in evaluation methodologies, limiting the reproducibility of experimental results. The adoption of common frameworks and standardized data pre-processing techniques significantly improves model transferability across domains. Through work of Demarchi et al. [5] the point about the lack of a standardized format leading to limiting progress is made.
Using a standardized implementation template, users can contribute new architectures, ensuring consistency while promoting collaboration. This design allows the dataset to evolve with advancements in machine learning. Transfer learning frameworks, such as those described in OpenTOS [23], have shown that integrating standardized templates can significantly accelerate model adaptation and optimization.
That leads us to extensibility, which ensures that neural network datasets remain adaptable as new architectures and training methodologies emerge. As proposed in [9] and through [30], a robust dataset should support user contributions, allowing researchers and developers to integrate novel models seamlessly. Furthermore, extensibility fosters collaborative development, which accelerates advancements in deep learning research. This approach ensures the framework to be more versatile and up to date with newest architectures. Additionally, we have introduced an accompanying SQLite-based database system, enabling efficient data management, query optimization, and streamlined integration with other components, to support the extensibility paradigm.
To align with our design principles, such a dataset should also be versatile enough to support multiple applications, including image classification, segmentation, and natural language processing, etc. Versatility enables the dataset to cater to various research domains. If a dataset is too narrow in scope, it restricts the range of architectures that can be developed. Considering the abundance of models, presented in [39], a neural network repository would be deemed as very limited.
Moving to the hyperparameter search possibility, our team decided on choosing Optuna [2]. It is an open-source framework designed for automated hyperparameter optimization. It employs techniques such as Tree-structured Parzen Estimator (TPE) sampling and multi-objective optimization to efficiently search for optimal hyperparameters. Unlike traditional grid or random search, Optuna dynamically explores promising hyperparameter configurations, significantly reducing computational overhead while improving model performance. Its use within many AutoML frameworks, validates its role in automated optimization [19].
The integration of statistical insights into the training process played a pivotal role in evaluating model performance. Accuracy was employed as the primary metric, with visualizations illustrating its progression over time. These visual representations provide valuable insights into the learning dynamics, indicating whether the model converges, diverges, or exhibits fluctuations in performance throughout the training process. From the reviewed neural network frameworks, this has also been done in [42, 31, 26] and more extensively in [40].
By integrating extensibility, standardization, versatility, advanced hyperparameter optimization, and statistical analysis, this work presents a comprehensive dataset of rigorously tested PyTorch neural network implementations and an advanced evaluation framework. The dataset addresses the demand for high-quality, standardized neural network implementations, offering a robust foundation for research and industry applications. It includes diverse architectures tailored for tasks like image classification and segmentation, with each model verified for correctness, stability, and performance. The novelty of LEMUR and its design, aimed at seamless AutoML use cases, makes it one of a kind. This uniqueness renders direct comparisons with other projects and repositories both impossible and obsolete, as LEMUR is specifically structured to provide a unified and automated benchmarking framework that existing resources lack. This work addresses current challenges in neural network design and analysis while paving the way for advancements in automated machine learning and code generation. It is a unique versatile resource for educators, researchers, and practitioners alike, serving as an all-in-one package combining everything needed for a neural network dataset.
@@ 2 LEMUR Overview
In this section we discuss the structure of the neural network dataset. The dataset is ever growing and includes 136 model implementations with 52 unique model architectures. The proposed LEMUR framework offers comprehensive set of innovative features. It provides a robust mechanism for tracking experiments, meticulously saving all essential hyperparameters and results from each experiment as explained in Sec. 4.1. This capability not only enables detailed analysis but also ensures reproducibility, which is critical for advancing research.
The framework integrates data into comprehensive statistical summaries, offering clear and actionable insights into model performance. This data is conveniently output in Excel format, making it accessible to users of varying skill levels and facilitating straightforward analysis. Furthermore, the framework includes automated visualization of key statistics, enabling enhanced interpretability of results and allowing researchers to focus on deriving meaningful conclusions and insights.
Designed with flexibility as a core principle, the LEMUR framework also supports seamless integration of additional scripts, enabling it to cater to a wide range of scientific investigations in neural network training. Together, these features create a robust, user-friendly framework that meets the needs of both novice and advanced researchers, empowering them to analyze, optimize, and interpret neural network performance with greater efficiency and precision.
The LEMUR framework streamlines data management by seamlessly integrating datasets while ensuring integrity and consistency. It employs unique data tracking, storing all records in JSON format with assigned identifiers to prevent duplication. During model training, the framework automatically logs configurations, metrics, and results into both JSON and a database, capturing task details, dataset identifiers, hyperparameters, and key performance metrics. This automation eliminates manual data handling, ensuring reliability and providing a structured foundation for analysis and visualization.
The general pipeline of the framework is depicted in Fig. 1. To initiate an experiment, the task, the neural networks, the dataloader for the task specific dataset, the transformation to be used in the preprocessing and the metric used for the training must be specified as shown in LABEL:listing:configuration_code.
The LEMUR pipeline begins by setting up the Optuna task dynamically according to the specified configuration. It then runs a number of Optuna trials with the goal of hyperparameter optimization, during which the results of each trial is recorded. These results are then saved as JSON files. To enable efficient data handling, the information from these JSON files is then recorded in a SQL database from which the graphical insights are automatically generated.
@@
python run.py -c <task-name>_<dataset-name>_<metric>_<neural-net-name>
This pipeline design ensures the flexibility, efficiency and extensibility that LEMUR is aiming for. Moreover, by storing experimental results in a structured database, the framework allows for seamless querying, comparison, and retrieval of model performance metrics across different configurations. The recorded data can be used to generate statistical summaries, performance trend analyses, and comparative visualizations, providing deeper insights into model behavior. Additionally, the modular nature of the pipeline enables users to extend the framework by integrating custom neural networks for arbitrary tasks, datasets, transformations, and evaluation metrics without modifying the core system.
@@ 4.1 Performance Exploration with Optuna
Optimizing model performance by identifying the best hyperparameters is a highly challenging task. Traditional approaches, such as manual grid search [3] or random search, are time-consuming and often inefficient. We have leveraged the efforts done in the field of bayesian hyperparameter optimization to incorporate model performance testing in our framework.
The baseline analysis employs Optuna [2], a state-of-the-art hyperparameter optimization framework, to systematically evaluate the performance of models within the LEMUR framework. Optuna is designed with a highly flexible and lightweight architecture, enabling seamless integration with a wide variety of machine learning workflows.
Optuna leverages the Tree-structured Parzen Estimator (TPE) [4], a Bayesian optimization method that efficiently explores hyperparameter spaces by adapting sampling based on prior evaluations. Unlike grid or random search, TPE focuses on promising regions, reducing computational overhead and accelerating convergence. This makes Optuna ideal for neural network optimization, particularly within the LEMUR framework, where it enables scalable benchmarking, automated trial management, and comprehensive statistical analysis of hyperparameter dynamics.
A unique strength of this implementation is its ability to handle hyperparameter spaces of arbitrary complexity, allowing users to explore any combination of parameters. This flexibility means that users can define as many hyperparameters as needed, tailoring optimization to the specific requirements of their neural network models.
Optuna automates experiment execution and statistical tracking, conducting iterative trials on key hyperparameters like learning rate, batch size, and momentum while ensuring consistency and eliminating manual intervention. Its built-in statistical logging captures metrics such as training accuracy, validation loss, and epoch-wise performance, streamlining experimentation and enabling comprehensive analysis. This automation enhances reproducibility, scalability, and benchmarking for large-scale datasets and diverse neural network architectures. Additionally, Optuna’s flexibility in parameter selection allows for customized experimentation, making it a powerful tool for fine-tuning and model comparison. By default, the number of recorded Optuna trials is 41,131, matching the size of the initial dataset.
The detailed experimental setups used for Optuna are provided in the supplementary material.
@@ 5 Conclusion & Future Work
In conclusion, this work introduces a multifaceted contribution to the field of machine learning by presenting a high-quality dataset of rigorously validated PyTorch neural network implementations and an advanced evaluation framework. The framework’s novelties—ranging from its extensibility with a standardized implementation template, to its integration with an SQLite-based database for efficient data management, and its robust tools for hyperparameter optimization and analysis—have been comprehensively explored. Together, these innovations address key challenges in neural network design and evaluation, enabling researchers and practitioners to benchmark, fine-tune, and extend architectures seamlessly. This resource not only supports diverse applications like AutoML, domain adaptation, and optimization research but also fosters collaboration and standardization, paving the way for future advancements in the field.
This framework is designed with growth and adaptability at its core, enabling the seamless addition of new neural network architectures and support for emerging tasks. By leveraging a standardized implementation template and a modular design, the framework ensures that contributions maintain consistency and interoperability, fostering a cohesive resource for researchers and practitioners. The extensibility of the framework allows it to evolve alongside advancements in machine learning, accommodating novel architectures and methodologies that address new challenges in diverse domains.
We invite the scientific community to collaborate and expand this resource by contributing new neural network implementations, datasets, and task-specific adaptations. By working together, we can grow the framework into a comprehensive and continuously updated repository that reflects the state-of-the-art in neural network research. Through such collaboration, the framework has the potential to broaden its applicability, support an even wider range of tasks, and drive innovation in areas such as automated machine learning, domain-specific transfer learning, and advanced optimization techniques. Let us collectively advance the horizons of neural network design and analysis by building upon this foundation.
We look forward to collaborating with the broader research community to drive innovation and enhance the capabilities of the neural network dataset. Your expertise and insights can make a significant impact on the future of this project!
Acknowledgment
We would like to express our gratitude towards Xi Zhang for his essential contributions in incorporating Bayesian and complex-valued neural networks, significantly enhancing the variety of deep learning model types.

### v4

@@ Abstract
Neural networks have become the backbone of modern AI, yet designing, evaluating, and comparing them remains labor-intensive. While many datasets exist for training models, there are few standardized collections of the models themselves. We present LEMUR, an open-source dataset and framework that brings together a large collection of PyTorch-based neural networks across tasks such as classification, segmentation, detection, and natural language processing. Each model follows a common template, with configurations and results logged in a structured database to ensure consistency and reproducibility. LEMUR integrates Optuna for automated hyperparameter optimization, provides statistical analysis and visualization tools, and exposes an API for seamless access to performance data. The framework also supports extensibility, enabling researchers to add new models, datasets, or metrics without breaking compatibility. By standardizing implementations and unifying evaluation, LEMUR aims to accelerate AutoML research, facilitate fair benchmarking, and lower the barrier to large-scale neural network experimentation. To encourage adoption and collaboration, LEMUR and its plugins are released as open-source projects under the MIT license and are available at: https://github.com/ABrain-One/nn-dataset, https://github.com/ABrain-One/nn-plots, and https://github.com/ABrain-One/nn-vr.
@@ 1 Introduction
The rapid progress of artificial intelligence is driven by increasingly powerful neural network architectures. Yet, designing, tuning, and benchmarking these models is still a slow and resource-heavy process that depends on extensive manual effort. Traditional datasets provide raw inputs such as images or text, but researchers lack standardized datasets of the models themselves—datasets that could make evaluation and automation far more efficient.
@@
With built-in support for image classification, segmentation, object detection, and natural language processing, LEMUR is a comprehensive resource for research and practical applications. It excels in AutoML by streamlining model selection and hyperparameter tuning while also supporting edge devices for efficient deployment in resource-constrained environments. Large language models, driven by their growing complexity and scale, are increasingly replacing traditional methods across various domains [6, 25], including those focused on the automated exploration of extensive hyperparameter spaces.
@@ 1.1 Related Work
In the domain of neural networks, datasets have always played a central role in enabling training, evaluation, and benchmarking. Large-scale labeled datasets such as ImageNet for vision tasks [26] and GLUE for natural language processing [29] have been the driving force behind many breakthroughs in deep learning. Over the years, a wide variety of resources have been developed for images, text, and even biological data [17, 13, 15, 28, 18]. These collections provide the raw inputs needed for learning representations. However, what has been missing is a curated dataset that captures the neural network architectures themselves—their structural variations, hyperparameters, and performance behaviors—rather than just the raw data they consume.
The motivation for building such a dataset grew out of the limitations of existing model libraries. For example, the PyTorch Vision Library [23] provides access to many popular pretrained models. Yet these models are not organized into a unified, standardized dataset that records architecture details, hyperparameters, and benchmarking statistics. Similarly, Hugging Face Transformers [31] has become a dominant hub for pretrained models in NLP and beyond, but, like PyTorch Vision, it does not enforce normalization or standardization that would allow seamless cross-domain evaluation or consistent workflows. In both cases, the models exist, but the infrastructure to treat them as part of a structured dataset is absent.
Another important context is the rise of AutoML frameworks, which aim to automate the optimization and evaluation of machine learning models. These frameworks have shown how standardized benchmarking and workflow normalization can accelerate progress and improve reproducibility [11, 19]. In parallel, advances in transfer learning [33] and novel loss functions [16] demonstrate the importance of having well-defined evaluation pipelines that make it easy to adapt, extend, and compare models. Despite this progress, no existing framework integrates these principles into a cohesive dataset of neural networks themselves. Current approaches automate training and tuning but do not address the lack of a structured, maintainable collection of models and their associated metadata.
Several attempts have been made to build repositories of neural networks for specific purposes. The CrossedWires dataset [34] highlights cross-framework reproducibility issues by showing that identical models implemented in different frameworks can produce significantly different accuracies. The Younger dataset [32] takes a different approach by compiling more than 7,600 unique AI-generated architectures from a pool of over 174,000 real-world models across 30 tasks. This dataset serves as a benchmark for graph neural networks and as a resource for studying architecture diversity. Other repositories fill educational or practical niches, such as Raschka’s deeplearning-models collection [20], which provides TensorFlow and PyTorch implementations mainly for teaching, or the Wolfram Neural Net Repository [24], which hosts pretrained models for immediate use. The ONNX Model Zoo [21] also contributes by enabling cross-framework compatibility with standardized ONNX models. Collectively, these resources enrich the ecosystem, but each is limited in scope: some focus on pretrained weights, others on architecture diversity, and others on framework compatibility.
The LEMUR dataset is designed to address this gap. It provides an expandable and standardized repository of PyTorch-based neural network implementations, each following a consistent template. Unlike ad-hoc collections, LEMUR couples these implementations with a unified evaluation and benchmarking framework, integrating hyperparameter tuning, structured logging, and automated statistical analysis. By treating models themselves as the “data,” LEMUR automates performance verification across training setups and data transformations, producing consistent performance statistics for each model. This design ensures that model evaluations are reproducible, comparable, and extensible.
We argue that LEMUR’s explicit focus on standardization, reproducibility, and unified evaluation positions it as a valuable addition to the deep learning ecosystem. It complements existing datasets and repositories by providing the missing layer of structured model-level benchmarking. This enables researchers not only to access diverse architectures but also to analyze them under a consistent framework, which is essential for progress in areas such as AutoML, model comparison, and reproducible research.
@@ 1.2 Our Contribution
LEMUR provides a unified infrastructure for benchmarking and extending PyTorch-based neural network models through a combination of standardized implementation templates, schema-driven configuration, and reproducible evaluation workflows. It supports a growing collection of diverse architectures for tasks such as classification and segmentation, each paired with YAML-based specs, automated hyperparameter tuning via Optuna, and detailed accuracy traces stored in an SQLite-backed database. All models are validated for correctness and convergence, with training statistics visualized to enable comparative analysis. By integrating these elements, LEMUR enables reliable experimentation, modular reuse, and seamless integration into AutoML frameworks.
@@ 2 LEMUR Overview
In this section, we discuss the structure of the neural network dataset and its supporting framework. The dataset is continuously expanding and currently includes 136 model implementations covering 52 unique architectures. The proposed LEMUR framework offers a comprehensive set of features designed to support reproducibility, extensibility, and analysis.
A central capability of LEMUR is its experiment tracking system. For every run, the framework automatically records all relevant configurations, including hyperparameters, dataset identifiers, and evaluation metrics, as detailed in Sec. 4.1. These records are stored in both JSON files and a SQL database, ensuring that no experiment is lost and that results can be easily retrieved or extended. This mechanism enables detailed analysis and guarantees reproducibility, which is critical for advancing neural network research.
To support interpretability, the framework integrates results into statistical summaries and automatically generates visualizations of key trends. In addition, all summaries can be exported in Excel format, making performance reports accessible to users with different technical backgrounds. This reduces the need for manual data handling and allows researchers to move directly from training to analysis and interpretation.
Flexibility is another defining principle of LEMUR. The framework supports the seamless integration of additional scripts, datasets, or metrics, making it suitable for a wide range of research scenarios. Its modular design ensures that both novice and advanced users can benefit: newcomers can run experiments and analyze results with minimal setup, while experienced researchers can extend the system to meet domain-specific needs.
The framework also streamlines data management. Unique identifiers are assigned to every record to prevent duplication, and all configurations, metrics, and results are logged automatically during training. By combining JSON storage for transparency with a database backend for structured queries, the framework creates a reliable foundation for large-scale benchmarking and reproducible research.
The general pipeline is depicted in Fig. 1. To launch an experiment, users specify the task, dataset loader, preprocessing transformations, evaluation metric, and neural network architecture, as illustrated in LABEL:listing:configuration_code.
@@
python run.py -c <task-name><dataset-name><metric>_<neural-net-name>
Once initialized, the pipeline sets up an Optuna optimization task dynamically according to the configuration. A series of trials is then executed, with each trial exploring different hyperparameter combinations. Results from each trial are first saved as JSON files and then transferred into the SQL database, from which graphical insights and comparative analyses are generated automatically.
This pipeline design ensures flexibility, efficiency, and extensibility. By storing results in a structured database, the framework allows seamless querying, comparison, and retrieval of performance metrics across models and tasks. These records can be used to generate statistical summaries, performance trend analyses, and comparative visualizations, providing deeper insights into model behavior. Moreover, the modular nature of the pipeline allows users to extend the framework with new architectures, tasks, transformations, and metrics without modifying the core system.
@@ 4.1 Performance Exploration with Optuna
Optimizing model performance by selecting effective hyperparameters is a persistent challenge. Traditional methods like grid or random search [3] are inefficient and computationally expensive. To address this, we incorporate Optuna [2], a modern hyperparameter optimization framework, into LEMUR for large-scale performance tuning.
Optuna employs the Tree-structured Parzen Estimator (TPE) [4], a Bayesian approach that adaptively focuses on promising regions of the hyperparameter space. This improves convergence and reduces compute costs. Its lightweight design integrates easily into modern ML workflows.
A key strength is its ability to handle complex, high-dimensional search spaces. Users can define arbitrary hyperparameter combinations tailored to their models. Optuna then automates the trial execution, testing parameters like learning rate, batch size, and momentum across diverse architectures.
Each trial is logged with detailed metrics—training accuracy, validation loss, and epoch-wise trends, allowing for thorough performance analysis. By default, 41,131 Optuna trials were conducted to match the size of the dataset.
This setup supports reproducible benchmarking and accelerates architecture-level comparisons. Optuna’s flexibility and automation make it well suited for scalable experimentation and integration with AutoML frameworks.
Full details on experimental setups and optimization configurations are provided in the supplementary material.
@@ 5 Conclusion & Future Work
In conclusion, this work introduces a comprehensive contribution to the machine learning community through the development of LEMUR, a high-quality dataset of rigorously validated PyTorch neural network implementations, paired with a unified evaluation and analysis framework. The framework combines several novel features: a standardized implementation template that ensures consistency across models, an SQLite-based database for structured experiment tracking and efficient data management, and integrated tools for hyperparameter optimization and statistical analysis Sec. 4.1. Together, these components address persistent challenges in neural network research, particularly those related to reproducibility, comparability, and extensibility. By providing a foundation where models can be benchmarked, fine-tuned, and extended under consistent conditions, LEMUR supports both fundamental research and practical applications in areas such as AutoML, domain adaptation, and optimization studies.
A defining aspect of the framework is its capacity for growth and adaptability. Built around modular principles, LEMUR allows new architectures, datasets, and task-specific methodologies to be added without breaking compatibility. This ensures that the framework can evolve alongside advances in deep learning, accommodating novel architectures, training strategies, and evaluation metrics as they emerge. The use of a standardized template also guarantees interoperability, meaning that future contributions integrate smoothly into the existing system while preserving reproducibility and consistency across the repository.
Looking forward, the framework is intended to serve not only as a resource but also as a collaborative platform for the community. By encouraging community contributions, LEMUR can expand into a comprehensive and continuously updated repository that reflects the state of the art in neural network design and experimentation. Such collaboration would broaden its applicability to new domains, support emerging tasks, and enhance its role in advancing reproducibility and standardized evaluation. In particular, extensions into areas such as automated machine learning, transfer learning for domain-specific applications, and advanced optimization strategies could be accelerated through shared development and use, ensuring that LEMUR remains relevant as the field evolves and adaptable to challenges that have yet to emerge.
Ultimately, we envision LEMUR as more than a static dataset or framework—it is designed as a foundation for collective progress. By offering consistency, transparency, and extensibility, it provides researchers and practitioners with the tools to not only conduct reproducible experiments but also to explore innovative approaches in a structured and reliable environment. As the framework grows, it has the potential to become a cornerstone resource for neural network research, lowering barriers for new investigations while supporting advanced studies in optimization, architecture design, and automated machine learning.